[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Feb 08, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Feb 8 15:01:13 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
DM sent the 18Z map on Sheephead. The smaller domain maps didn't update on  
the Web. They went out okay on Cumulus. Matt Sardi was not able to  
replicate the error, but modified the script to add to the log file to help  
with troubleshooting if the problem happens again.

The OPC backup yesterday went a little longer than expected, but was  
completed by early evening. Thanks to Andy Latto and Chris Mello for  
staying late and seeing this through to the end.
HSU:
NCO provided some new instructions for the tuna N-AWIPS workstation under  
RHEL6. Matt will be working with Brian to continue the testing.
Other items:
Phil Cragg from NCO restored the local GOES ingest systems (east and west).  
He will also provide documentation so we can more easily troubleshoot the  
ingest from NHC.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170208/87430d7b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list